# Multi-scale tests of independence powerful for detecting explicit or implicit functional relationship

### v1

@@ Abstract
In this article, we consider the problem of testing the independence between two random variables. Our primary objective is to develop tests that are highly effective at detecting associations arising from explicit or implicit functional relationship between two variables. We adopt a multi-scale approach by analyzing neighborhoods of varying sizes within the dataset and aggregating the results. We introduce a general testing framework designed to enhance the power of existing independence tests to achieve our objective. Additionally, we propose a novel test method that is powerful as well as computationally efficient. The performance of these tests is compared with existing methods using various simulated datasets.
@@ 1 Introduction
The article is organized as follows. In Section 2, we introduce a general testing framework aimed at enhancing the power of existing tests of independence in scenarios described by Equation (2) and examine its computational complexity. In Section 3, we apply this framework to develop a novel test of independence and analyze its computational complexity. Section 4 presents a performance comparison of our methods on various simulated datasets. Finally, we conclude with a discussion and summary of our findings.
2 A General Testing Framework
@@
  • Example C: $Y=U+\epsilon_{Y}$ and $X=U^{2}+\epsilon_{X}$, where $U\sim U(0,1)$ and $U$ is independent of $\epsilon_{X}$ and $\epsilon_{Y}$.
@@
As the next example, we consider the $BEX_{d}$ distribution, originally introduced in (Zhang, 2019) where it is defined as the uniform distribution over a set of parallel and intersecting lines given by $\sum_{i=1}^{2^{d}-1}\sum_{j=1}^{2^{d}-1}(|x-c_{i}|-|y-c_{j}|)I[|x-c_{i}|\leq 2^{-d},|y-c_{j}|\leq 2^{-d}]=0$, where $c_{i}=(2i-1)/2^{d}$. Figure 2 illustrates the $BEX_{d}$ distribution for $d=2,3$, and $4$. If $(X,Y)$ follows the $BEX_{d}$ distribution, the marginals $X$ and $Y$ are dependent but both follow a continuous uniform distribution over $[0,1]$. Moreover, $X$ and $Y$ can be shown to satisfy Equation 2. Although $\text{Cor}(X,Y)=0$ for the $BEX_{d}$ distribution, there exist neighborhoods around support points where the conditional correlation is extremely high.
@@
These examples suggest that by calculating the test statistics for an existing test of independence across different neighborhoods and aggregating these values meaningfully, we can enhance its power, especially in scenarios described by Equation (2). Building on this idea, we introduce a multi-scale approach for existing tests of independence in this section.
@@
For each observation $(x_{i},y_{i})$, we order remaining observations according to their Euclidean distance from $(x_{i},y_{i})$. Random tie-breaking is to be used in case of ties while ordering. Let $(x_{\pi_{i}(1)},y_{\pi_{i}(1)}),$ $\ldots$,… , $(x_{\pi_{i}(n-1)},y_{\pi_{i}(n-1)})$ be the observations ordered by their Euclidean distance from $(x_{i},y_{i})$ in ascending order. Thus, $N_{i,\pi_{i}(k)}$ is the neighborhood of $(x_{i},y_{i})$ that has $k$-th nearest neighbor of $(x_{i},y_{i})$ at one of its vertices. It is easy to see that for a fixed $i$, the length of the diagonal of $N_{i,\pi_{i}(k)}$ increases as $k$ increases. We average the value of the test statistic $T$ based on the sample $\mathbf{xy}_{1:n}\cap N_{i,\pi_{i}(k)}$ keeping $k$ fixed and varying over $i$ over $1,\ldots,n$. We denote this average as $T_{[k]}$, that is, $T_{[k]}:=n^{-1}\sum_{i=1}^{n}T(\mathbf{xy}_{1:n}\cap N_{i,\pi_{i}(k)})=n^{-1}\sum_{i=1}^{n}T_{i,\pi_{i}(k)}$. To determine how extreme the value of $T_{[k]}$ is, we need to know its distribution under the independence of the marginals, which can be estimated by resampling technique. We discuss this in details next.
@@
In this illustration, we consider two popular test statistics as $T$: absolute value of Pearson’s correlation and distance correlation. We denote these two test statistics by $T^{cor}$ and $T^{dcor}$ respectively. We consider eight bivariate distributions, description of each is provided in the Table 1.
@@
(h) BVN, rho=0.5 $(X,Y)$ follows bivariate normal distribution
with zero means, unit variances and correlation
coefficient 0.5
@@
From Figure 4, we observe that the average Z-score values are close to 0 for ‘Square’ distribution as $X$ and $Y$ are independent. For the ‘Straight line’ and ‘Noisy straight line’ distributions, the average Z-scores are higher in larger neighborhoods for both $T^{cor}$ and $T^{dcor}$. However, in the ‘Noisy straight line’, the dependence information is less apparent in smaller neighborhoods compared to the ‘Straight line’. For the ‘Sine’ distribution, the dependence information is clearly noticeable in smaller neighborhoods. An interesting phenomenon occurs with the ‘Circle’ distribution, where most neighborhoods, including the largest ones, contain arcs, allowing both $T^{cor}$ and $T^{dcor}$ to detect dependence effectively in most neighborhoods. In the ‘Noisy parabola’, the dependence is most prominent in mid-sized neighborhoods. For the more complex $BEX_{2}$ distribution, these statistics detect the dependence patterns in smaller and mid-sized neighborhoods, but less so in larger ones. Finally, in ‘BVN, rho=0.5’, the dependence information is only clearly evident in larger neighborhoods.
@@
First, we observe that under independence, the expected value of $z_{k}$ is 00 for $k=1,\ldots,n-1$. Under dependence, if the $T_{i,j}$’s are stochastically larger, $T_{[k]}$ will also be stochastically larger. In that case, the expected values of $z_{k}$ will also be positive. Let $\mu_{k}$ be the expected value of $z_{k}$, that is, $\mu_{k}=E[z_{k}]$. Thus testing $H_{0}$ against $H_{1}$ can be done by testing $H_{0}^{{}^{\prime}}:\mu_{k}=0~{}\forall k$ vs. $H_{1}^{{}^{\prime}}:\exists~{}k~{}\text{such that}~{}\mu_{k}>0$. To test this, we suggest the following test statistic $\Psi_{n}:=\sum_{k=1}^{n-1}(\max\{z_{k},0\})^{2}$. Clearly, a high value of $\Psi_{n}$ presents evidence against null hypothesis. To determine the distribution of $\Psi_{n}$ under $H_{0}$, we again use resampling approach by utilizing $B$ randomly permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{B})}$ which are already in our disposal. Similar how we computed $z_{1},\ldots,z_{n-1}$ for $\mathbf{xy}_{1:n}$ using permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{B})}$, we compute $z_{1}^{(\tau_{i})},\ldots,z_{n-1}^{(\tau_{i})}$ for $\mathbf{xy}_{1:n}^{(\tau_{i})}$ based on permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{i-1})},\mathbf{xy}_{1:n}^{(\tau_{i+1})},\mathbf{xy}_{1:n}^{(\tau_{B})}$ for $i=1,\ldots,B$. Next we calculate $\Psi_{n}^{(\tau_{i})}:=\sum_{k=1}^{n-1}(\max\{z_{k}^{(\tau_{i})},0\})^{2}$ for $i=1,\ldots,B$. Finally we determine the p-value of $\Psi_{n}$ by $p=B^{-1}\sum_{i=1}^{B}I[\Psi_{n}\leq\Psi_{n}^{(\tau_{i})}]$. We reject $H_{0}$ if the p-value turns out to be less than the level of significance.
@@ 3 A Special Test Method
As shown in Figure 5(a), the Z-scores for $T^{phi}$ exhibit higher variance compared to $T^{cor}$ under independence. Figures 5(b) through 5(g) demonstrate that $T^{phi}$ consistently yields higher Z-scores than $T^{cor}$ for the ‘Straight line,’ ‘Noisy straight line,’ ‘Sine,’ ‘Circle,’ ‘noisy parabola,’ and ‘$BEX_{2}$’ distributions. Additionally, the average Z-score plot for $T^{phi}$ appears to be more wiggly compared to $T^{cor}$. In the ‘BVN, rho=0.5’ distribution, the average Z-scores of $T^{phi}$ are comparable to those of $T^{cor}$.
@@ 4 Performance on simulated datasets
In this section, we compare the performance of our proposed test methods on various simulated datasets with that of existing tests in the literature. From the proposed general testing framework, we select two well-known test statistics - the absolute value of Pearson’s correlation and distance correlation - as $T$. We denote the resulting test statistic $\Psi_{n}$ in these two cases as $\Psi_{n}^{cor}$ and $\Psi_{n}^{dcor}$, respectively. We denote the test statistic $\Psi_{n}$ for our proposed special test method as $\Psi_{n}^{phi}$. As pointed out in previous sections, the computational complexities of $\Psi_{n}^{phi},\Psi_{n}^{cor}$, and $\Psi_{n}^{dcor}$ are $\mathcal{O}(n^{2}\log n),\mathcal{O}(n^{3})$, and $\mathcal{O}(n^{3}\log n)$, respectively.
From existing tests of independence, we selected the following popular tests: dCor (Székely et al., 2007), HSIC (Gretton et al., 2007), HHG (Heller et al., 2013), MIC (Reshef et al., 2011), Chatterjee’s correlation (xicor) (Chatterjee, 2021) and BET (Zhang, 2019). For performing dCor, HSIC, HHG, and xicor tests, we used R packages energy (Rizzo and Szekely, 2022), dHSIC (Pfister and Peters, 2019), HHG (Kaufman et al., 2019), and XICOR (Chatterjee and Holmes, 2023), respectively. We calculated the MIC statistic by using ‘mine_stat’ function from the R package minerva (Albanese et al., 2012) and subsequently performed a permutation test. For executing BET test, we used the R code provided in the supplemental material of (Zhang, 2019). We fixed the nominal level at $5\%$. Except for BET, all other tests were performed using the permutation principle, with the number of permutations set to 1000. The empirical power of a test is determined by calculating the proportion of times it rejects the null hypothesis over 1000 independently generated samples of a specific size from a given distribution.
First, we considered eighty jointly continuous distributions, namely ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous A’, ‘Lissajous B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’. Their descriptions are provided in Table 3. Scatter plots for each of these distributions are presented in Figure 8 in Appendix B.
@@
(c) Lissajous A $\Theta\sim U(0,2\pi),X=\sin(3\Theta+3\pi/4)+\epsilon_{X},Y=\sin(\Theta)+\epsilon_{Y}$
(d) Lissajous B $\Theta\sim U(0,2\pi),X=\sin(4\Theta+\pi/2)+\epsilon_{X},Y=\sin(3\Theta)+\epsilon_{Y}$
@@
Table 3: The description of ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous A’, ‘Lissajous B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’ distributions. Here $\epsilon_{X},\epsilon_{Y}$ are i.i.d. random variables following $N(0,1/30^{2})$, independent of $U,V$, and $\Theta$.
@@
The empirical powers of all test methods on these 8 distributions are presented in Figure 6.
Figure 6: The empirical powers of different test methods on ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous A’, ‘Lissajous B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’.
It can be seen from this figure that $\Psi_{n}^{phi}$ performed best in both ‘Doppler’ and ‘$BEX_{2}$’ and performed well in all other distributions. $\Psi_{n}^{cor}$ performed fairly well in all distributions except for ‘Rose curve’ and ‘Tilted square’. $\Psi_{n}^{dcor}$ performed well in all cases, and in ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Five clouds’, it performed the best. HHG yielded satisfactory power overall except for ‘Lissajous curve B’. The power of HSIC is satisfactory only in the ‘Doppler’ and ‘Tilted square’ distributions, not otherwise. dCor, MIC, xicor did not perform well. BET performed somewhat well for larger sample sizes in ‘Lissajous curve B’, ‘Rose curve’ and in ‘Spiral’.
Next, we considered four datasets to assess the performance of our methods under different levels of noise. In these examples, $X=f(Z)+\lambda\epsilon_{X}$ and $Y=g(Z)+\lambda\epsilon_{Y}$, where $f$ and $g$ are real functions, $Z$ is a random variable, $\lambda$ is a non-negative constant, and $\epsilon_{X}$ and $\epsilon_{Y}$ are i.i.d. normal random variables that are also independent of $Z$. As $\lambda$ increases, the noise level increases. In particular, we considered 4 distribution named - ‘Parabola $\lambda$’, ‘Circle $\lambda$’, ‘Sine $\lambda$’ and ‘Laminscate $\lambda$’, descriptions of each of these can be found in Table 4.

### v2

# A Novel Multiscale Framework for Testing Independence: Efficient Detection of Explicit or Implicit Functional Relationships

@@ Abstract
In this article, we consider the problem of testing the independence between two random variables. Our primary objective is to develop tests that are highly effective at detecting associations arising from explicit or implicit functional relationship between two variables. We adopt a multiscale approach by analyzing neighborhoods of varying sizes within the dataset and aggregating the results. We introduce a general testing framework designed to enhance the power of existing independence tests to achieve our objective. Additionally, we propose a novel test method that is powerful as well as computationally efficient. The performance of these tests is compared with existing methods using various simulated datasets. Additionally, a visualization method has been proposed for exploring the localization of dependence within datasets.
@@ 1 Introduction
The article is organized as follows. In Section 2, we introduce a general testing framework aimed at enhancing the power of existing tests of independence in scenarios described by Equation (2) and examine its computational complexity. In Section 3, we apply this framework to develop a novel test of independence and analyze its computational complexity. Section 4 presents a performance comparison of our methods on various simulated datasets. Section 5 extends this analysis to a real-world dataset. In Section 6, we demonstrate a visualization technique for exploring the localization of dependence in a real-world dataset. Finally, we conclude with a discussion and summary of our findings.
2 A General Multiscale Testing Framework
@@
  • Example C: $Y=U+\epsilon_{Y}$ and $X=U^{2}+\epsilon_{X}$, where $U\sim U(-1,1)$ and $U$ is independent of $\epsilon_{X}$ and $\epsilon_{Y}$.
@@
As the next example, we consider the $BEX_{d}$ distribution, originally introduced in (Zhang, 2019) where it is defined as the uniform distribution over a set of parallel and intersecting lines given by $\sum_{i=1}^{2^{d}-1}\sum_{j=1}^{2^{d}-1}(|x-c_{i}|-|y-c_{j}|)I[|x-c_{i}|\leq 2^{-d},|y-c_{j}|\leq 2^{-d}]=0$, where $c_{i}=(2i-1)/2^{d}$. Figure 2 illustrates the $BEX_{d}$ distribution for $d=1,2$, and $3$. If $(X,Y)$ follows the $BEX_{d}$ distribution, the marginals $X$ and $Y$ are dependent but both follow a continuous uniform distribution over $[0,1]$. Moreover, $X$ and $Y$ can be shown to satisfy Equation 2. Although $\text{Cor}(X,Y)=0$ for the $BEX_{d}$ distribution, there exist neighborhoods around support points where the conditional correlation is extremely high.
@@
These examples suggest that by calculating the test statistics for an existing test of independence across different neighborhoods and aggregating these values meaningfully, we can enhance its power, especially in scenarios described by Equation (2). Building on this idea, we introduce a multiscale approach for existing tests of independence in this section.
@@
For each observation $(x_{i},y_{i})$, we order the remaining observations according to their Euclidean distance from $(x_{i},y_{i})$. Random tie-breaking is to be used in case of ties while ordering. Let $(x_{\pi_{i}(1)},y_{\pi_{i}(1)}),$ $\ldots$,… , $(x_{\pi_{i}(n-1)},y_{\pi_{i}(n-1)})$ be the observations ordered by their Euclidean distance from $(x_{i},y_{i})$ in ascending order. Thus, $N_{i,\pi_{i}(k)}$ is the neighborhood of $(x_{i},y_{i})$ that has $k$-th nearest neighbor of $(x_{i},y_{i})$ at one of its vertices. It is easy to see that for a fixed $i$, the length of the diagonal of $N_{i,\pi_{i}(k)}$ increases as $k$ increases. We average the value of the test statistic $T$ based on the sample $\mathbf{xy}_{1:n}\cap N_{i,\pi_{i}(k)}$ keeping $k$ fixed and varying over $i$ over $1,\ldots,n$. We denote this average as $T_{[k]}$, that is, $T_{[k]}:=n^{-1}\sum_{i=1}^{n}T(\mathbf{xy}_{1:n}\cap N_{i,\pi_{i}(k)})=n^{-1}\sum_{i=1}^{n}T_{i,\pi_{i}(k)}$. To determine how extreme the value of $T_{[k]}$ is, we need to know its distribution under the independence of the marginals, which can be estimated by resampling technique. We discuss this in detail next.
Figure 3: Given the four observation points $(1,2)$, $(2,3)$, $(-1,-1)$, and $(5,4)$, the areas enclosed by the blue, black, and orange rectangles represent the neighborhoods selected around the observation point $(1,2)$.
@@
In this illustration, we consider two popular test statistics as $T$: absolute value of Pearson’s correlation and distance correlation. We denote these two test statistics by $T^{cor}$ and $T^{dcor}$ respectively. We consider eight bivariate distributions, and a description of each is provided in Table 1.
@@
(h) BVN, rho=0.5 $(X,Y)$ follows standard bivariate normal
distribution with a correlation coefficient of 0.5
@@
From Figure 5, we observe that the average Z-score values are close to 0 for ‘Square’ distribution as $X$ and $Y$ are independent. For the ‘Straight line’ and ‘Noisy straight line’ distributions, the average Z-scores are higher in larger neighborhoods for both $T^{cor}$ and $T^{dcor}$. However, in the ‘Noisy straight line’, the dependence information is less apparent in smaller neighborhoods compared to the ‘Straight line’. For the ‘Sine’ distribution, the dependence information is clearly noticeable in smaller neighborhoods. An interesting phenomenon occurs with the ‘Circle’ distribution, where most neighborhoods, including the largest ones, contain arcs, allowing both $T^{cor}$ and $T^{dcor}$ to detect dependence effectively in most neighborhoods. In the ‘Noisy parabola’, the dependence is most prominent in mid-sized neighborhoods. For the more complex ‘$BEX_{2}$’ distribution, these statistics detect the dependence patterns in smaller and mid-sized neighborhoods, but less so in larger ones. Finally, in ‘BVN, rho=0.5’, the dependence information is only clearly evident in larger neighborhoods.
@@
First, we observe that under independence, the expected value of $z_{k}$ is 00 for $k=1,\ldots,n-1$. Under dependence, if the $T_{i,j}$’s are stochastically larger, $T_{[k]}$ will also be stochastically larger. In that case, the expected values of $z_{k}$ will also be positive. Let $\mu_{k}$ be the expected value of $z_{k}$, that is, $\mu_{k}=E[z_{k}]$. Thus testing $H_{0}$ against $H_{1}$ can be done by testing $H_{0}^{{}^{\prime}}:\mu_{k}=0~{}\forall k$ vs. $H_{1}^{{}^{\prime}}:\exists~{}k~{}\text{such that}~{}\mu_{k}>0$. To test this, we suggest the following test statistic $\Psi_{n}:=\sum_{k=1}^{n-1}(\max\{z_{k},0\})^{2}$. Clearly, a high value of $\Psi_{n}$ presents evidence against the null hypothesis. To determine the distribution of $\Psi_{n}$ under $H_{0}$, we again use resampling approach by utilizing $B$ randomly permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{B})}$ which are already in our disposal. Similar how we computed $z_{1},\ldots,z_{n-1}$ for $\mathbf{xy}_{1:n}$ using permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{B})}$, we compute $z_{1}^{(\tau_{i})},\ldots,z_{n-1}^{(\tau_{i})}$ for $\mathbf{xy}_{1:n}^{(\tau_{i})}$ based on permuted samples $\mathbf{xy}_{1:n}^{(\tau_{1})},\ldots,\mathbf{xy}_{1:n}^{(\tau_{i-1})},\mathbf{xy}_{1:n}^{(\tau_{i+1})},\mathbf{xy}_{1:n}^{(\tau_{B})}$ for $i=1,\ldots,B$. Next we calculate $\Psi_{n}^{(\tau_{i})}:=\sum_{k=1}^{n-1}(\max\{z_{k}^{(\tau_{i})},0\})^{2}$ for $i=1,\ldots,B$. Finally we determine the p-value of $\Psi_{n}$ by $p=B^{-1}\sum_{i=1}^{B}I[\Psi_{n}\leq\Psi_{n}^{(\tau_{i})}]$. We reject $H_{0}$ if the p-value turns out to be less than the level of significance.
@@
Algorithm for General Multiscale Testing Framework Input: set of bivariate observations $xy_{1:n}$, test statistics $T$, number of permutations $B$ 1. Set $i=1$. 2. For $k=1,\dots,n-1$, determine $\pi_{i}(k)$ such that $(x_{\pi_{i}(k)},y_{\pi_{i}(k)})\in xy_{1:n}$ is the $k$-th nearest neighbor of $(x_{i},y_{i})$ in terms of Euclidean distance. Here $(x_{i},y_{i})$ is the $i$-th observation in $xy_{1:n}$. 3. If $i<n$, increase $i$ by 1 and go to Step 2. 4. For $k=1,\ldots,n-1$, compute $T_{[k]}:=\frac{1}{n}\sum_{i=1}^{n}T(xy_{1:n}\cap N_{i,\pi_{i}(k)}),$ where $N_{i,\pi_{i}(k)}:=[x_{i}-|x_{i}-x_{\pi_{i}(k)}|,x_{i}+|x_{i}-x_{\pi_{i}(k)}|]\times[y_{i}-|y_{i}-y_{\pi_{i}(k)}|,y_{i}+|y_{i}-y_{\pi_{i}(k)}|]$. 5. Set $b=1$. 6. Generate $xy^{(\tau_{b})}_{1:n}:=\{(x_{1},y_{\tau_{b}(1)}),\dots,(x_{n},y_{\tau_{b}(n)})\}$, where $\tau_{b}$ is a random permutation on $\{1,\dots,n\}$. Here $(x_{i},y_{i})$ is the $i$-th observation in $xy_{1:n}$. 7. Perform Steps 1 to 4 using $xy^{(\tau_{b})}_{1:n}$ instead of $xy_{1:n}$ to compute $T^{(\tau_{b})}_{[1]},\ldots,T^{(\tau_{b})}_{[n-1]}$. 8. If $b<B$, increase $b$ by 1 and go to Step 6. 9. Calculate $\bar{T}^{H_{0}}_{[k]}:=\frac{1}{B}\sum_{i=1}^{B}T^{(\tau_{i})}_{[k]}$ and $s^{H_{0}}_{[k]}:=\sqrt{\frac{1}{B-1}\sum_{i=1}^{B}\left(T^{(\tau_{i})}_{[k]}-\bar{T}^{H_{0}}_{[k]}\right)^{2}}$. 10. For $k=1,2,\dots,n-1$, compute $z_{k}:=\frac{T_{[k]}-\bar{T}^{H_{0}}_{[k]}}{s^{H_{0}}_{[k]}}$. Set $z_{k}:=0$ if $s^{H_{0}}_{[k]}=0$. 11. Compute $\Psi_{n}:=\sum_{k=1}^{n-1}\left(\max\{z_{k},0\}\right)^{2}$. 12. Reset $b=1$. 13. For $k=1,2,\dots,n-1$, compute $z_{k}^{(\tau_{b})}:=\frac{T^{(\tau_{b})}_{[k]}-\bar{T}^{H_{0}}_{[k]}}{s^{H_{0}}_{[k]}}$. Set $z_{k}^{(\tau_{b})}:=0$ if $s^{H_{0}}_{[k]}=0$. 14. Compute $\Psi_{n}^{(\tau_{b})}:=\sum_{k=1}^{n-1}\left(\max\{z_{k}^{(\tau_{b})},0\}\right)^{2}$. 15. If $b<B$, increase $b$ by 1 and go to Step 13. 16. Determine the p-value of $\Psi_{n}$ as: $p=\frac{1}{B}\sum_{i=1}^{B}I[\Psi_{n}\leq\Psi^{(\tau_{i})}_{n}]$.
@@ 3 A Special Test Method
As shown in Figure 6(a), the Z-scores for $T^{phi}$ exhibit higher variance compared to $T^{cor}$ under independence. Figures 6(b) through 6(g) demonstrate that $T^{phi}$ consistently yields higher Z-scores than $T^{cor}$ for the ‘Straight line,’ ‘Noisy straight line,’ ‘Sine,’ ‘Circle,’ ‘Noisy parabola,’ and ‘$BEX_{2}$’ distributions. Additionally, the average Z-score plot for $T^{phi}$ appears to be more wiggly compared to $T^{cor}$. In the ‘BVN, rho=0.5’ distribution, the average Z-scores of $T^{phi}$ are comparable to those of $T^{cor}$.
@@ 4 Performance on simulated datasets
In this section, we compare the performance of our proposed test methods on various simulated datasets with that of existing tests in the literature. From the proposed general testing framework, we select two well-known test statistics - the absolute value of Pearson’s correlation and distance correlation - as $T$. We denote the resulting test statistic $\Psi_{n}$ in these two cases as $\Psi_{n}^{cor}$ and $\Psi_{n}^{dcor}$, respectively. We denote the test statistic $\Psi_{n}$ for our proposed special test method as $\Psi_{n}^{phi}$. As pointed out in previous sections, the computational complexities of $\Psi_{n}^{phi},\Psi_{n}^{cor}$, and $\Psi_{n}^{dcor}$ are $\mathcal{O}(n^{2}\log n),\mathcal{O}(n^{3})$, and $\mathcal{O}(n^{3}\log n)$, respectively. For our proposed tests, we created an R package, ‘PAIT,’ which is available at https://github.com/angshumanroycode/PAIT.
From existing tests of independence, we selected the following popular tests: dCor (Székely et al., 2007), HSIC (Gretton et al., 2007), HHG (Heller et al., 2013), MIC (Reshef et al., 2011), Chatterjee’s correlation (xicor) (Chatterjee, 2021) and BET (Zhang, 2019). For performing dCor, HSIC, HHG, and xicor tests, we used R packages energy (Rizzo and Szekely, 2022), dHSIC (Pfister and Peters, 2019), HHG (Kaufman et al., 2019), and XICOR (Chatterjee and Holmes, 2023), respectively. We calculated the MIC statistic by using the ‘mine_stat’ function from the R package minerva (Albanese et al., 2012) and subsequently performed a permutation test. For executing BET test, we used the R code provided in the supplemental material of (Zhang, 2019). We fixed the nominal level at $5\%$. Except for BET, all other tests were performed using the permutation principle, with the number of permutations set to 1000. The empirical power of a test is determined by calculating the proportion of times it rejects the null hypothesis over 1000 independently generated samples of a specific size from a given distribution.
To ensure our proposed tests adhere to the specified significance level, we calculated the empirical power for two bivariate distributions with independent marginals. In the first case, both marginals followed standard normal distributions, while in the second case, they followed uniform distributions over $[0,1]$. The significance level was set at 5%. Table 3 presents a summary of the empirical power of various tests for sample sizes of 10, 20, and 50.
$n$ $\Psi_{n}^{phi}$ $\Psi_{n}^{cor}$ $\Psi_{n}^{dcor}$ HHG HSIC dCor MIC XICOR BET
Independent $N(0,1)$ marginals
10 0.048 0.058 0.059 0.041 0.051 0.053 0.008 0.037 0.003
20 0.047 0.046 0.050 0.042 0.050 0.034 0.046 0.040 0.034
50 0.044 0.050 0.041 0.055 0.055 0.053 0.060 0.061 0.053
Independent $U(0,1)$ marginals
10 0.051 0.047 0.047 0.052 0.048 0.045 0.009 0.032 0.000
20 0.050 0.049 0.050 0.041 0.040 0.054 0.044 0.037 0.040
50 0.052 0.053 0.053 0.049 0.048 0.057 0.055 0.050 0.038
Table 3: Empirical power of test methods under independent marginals.
All test methods, except MIC, XICOR, and BET, maintained an empirical power close to 5% across all sample sizes. MIC, XICOR, and BET exhibited conservative behavior for small sample sizes but approached the nominal significance level as the sample size increased.
First, we considered eight jointly continuous distributions, namely ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’. Their descriptions are provided in Table 4. Scatter plots for each of these distributions are presented in Figure 10 in Appendix B.
@@
(c) Lissajous curve A $\Theta\sim U(0,2\pi),X=\sin(3\Theta+3\pi/4)+\epsilon_{X},Y=\sin(\Theta)+\epsilon_{Y}$
(d) Lissajous curve B $\Theta\sim U(0,2\pi),X=\sin(4\Theta+\pi/2)+\epsilon_{X},Y=\sin(3\Theta)+\epsilon_{Y}$
@@
Table 4: The description of ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’ distributions. Here $\epsilon_{X},\epsilon_{Y}$ are i.i.d. random variables following $N(0,1/30^{2})$, independent of $U,V$, and $\Theta$.
@@
The empirical powers of all test methods on these eight distributions are presented in Figure 7.
Figure 7: The empirical powers of different test methods on ‘Doppler’, ‘$BEX_{2}$’, ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Tilted square’, and ‘Five clouds’ distributions.
It can be seen from this figure that $\Psi_{n}^{phi}$ performed best in both ‘Doppler’ and ‘$BEX_{2}$’ and performed well in all other distributions. $\Psi_{n}^{cor}$ performed fairly well in all distributions except for the ‘Rose curve’ and ‘Tilted square.’ $\Psi_{n}^{dcor}$ performed well in all cases, and in ‘Lissajous curve A’, ‘Lissajous curve B’, ‘Rose curve’, ‘Spiral’, ‘Five clouds’, it performed the best. HHG yielded satisfactory power overall except for ‘Lissajous curve B’. The power of HSIC is satisfactory only in the ‘Doppler’ and ‘Tilted square’ distributions, not otherwise. dCor, MIC, xicor did not perform well. BET performed well for larger sample sizes in ‘Lissajous curve B’, ‘Rose curve’ and in ‘Spiral’.
Next, we considered four datasets to assess the performance of our methods under different levels of noise. In these examples, $X=f(Z)+\lambda\epsilon_{X}$ and $Y=g(Z)+\lambda\epsilon_{Y}$, where $f$ and $g$ are continuous non-constant functions, $Z$ is a random variable, $\lambda$ is a non-negative constant, and $\epsilon_{X}$ and $\epsilon_{Y}$ are i.i.d. normal random variables that are also independent of $Z$. As $\lambda$ increases, the noise level increases. In particular, we considered 4 distribution named - ‘Parabola $\lambda$’, ‘Circle $\lambda$’, ‘Sine $\lambda$’ and ‘Laminscate $\lambda$’, descriptions of each of these can be found in Table 5.